OpenStudy (anonymous):

Simplify (44w + 2)(44w + 2)

OpenStudy (anonymous):

Solve for w over the real numbers: (2+44 w)^2 = 0 Eliminate the exponent. Take the square root of both sides: 2+44 w = 0 Factor the left hand side. Factor constant terms from the left hand side: 2 (1+22 w) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 2: 1+22 w = 0 Isolate terms with w to the left hand side. Subtract 1 from both sides: 22 w = -1 Solve for w. Divide both sides by 22: Answer: | | w = -1/22

OpenStudy (anonymous):

FOIL: First Outer Inner Last (44w + 2)(44w + 2) 1936w²+88w+88w+4 1936w²+176w+4 4(484w²+44w+1) (4)(484w²+22w+22w+1) (4) 22w(22w+1)+ 1(22w+1) 4(22w+1)(22w+1)